# Does synthetic data help low-resource Named Entity Recognition?

### v1

@@ Appendix A Related-Language Model Details
Appendix B LLM Prompts & Implementation
{m examples}
Appendix C Full Results of Training on Synthetic Data
C.1 NER From Scratch Setting
Figure 3 is a plot of results of training on synthetic data, in the ner from scratch setting, for all languages.
C.2 NER Fine-tuning Setting
Figure 4 is a plot of results of training on synthetic data, in the ner fine-tuning setting, for all languages.
Appendix D Full Results of WikiANN Comparison
Below in Table 5 is the full list of comparisons between NER model performance when trained on organic data, GPT-4-produced data, and WikiANN data.
@@ Appendix D Full Results of WikiANN Comparison
Figure 3: ner from scratch Setting
Figure 4: ner fine-tuning Setting

### v2

@@ Appendix B Related-Language Model Details
B.1 NER-fine tuning: Implementation Details
Appendix C Full Results of WikiANN Comparison
The WikiANN dataset is a massively multilingual NER benchmark, comprising data from 176 languages Pan et al. (2017); Rahimi et al. (2019). [^4] As Lignos et al. (2022) also note, strictly speaking, the original version of WikiANN put together by Pan et al. (2017) contains data from 282 languages; the version of the dataset commonly downloaded from Huggingface, however, and put together by Rahimi et al. (2019), contains data from 176 languages. In this work, we refer to the latter when referring to the WikiANN dataset. Table 3 shows the full list of comparisons between NER model performance when trained on organic data, GPT-4-produced data, and WikiANN data. The sizes of the WikiANN train sets vary significantly between different languages, meaning we often cannot assess the quality of the data in the context of training sets containing over 1000 datapoints (e.g. Kannada and Yoruba, whose WikiANN train sets contain only 100 datapoints). In such cases, however, we compare model performance when trained on equally small amounts of organic or LLM-produced synthetic data.